# Mapping fNIRS Signals to Agent Performance: Toward Reinforcement Learning from Neural Feedback

### v1

@@ Results
Single-Subject Models: Single-subject models were able to distinguish between both binary ($F1=0.79$) and multi-class ($F1=0.75$) levels of agent performance. Multi-class classifiers were able to differentiate between optimal, sub-optimal, and worst-case behavior with little variance. However, cross-validation studies revealed that no single-subject model was able to predict agent performance from the neural data of another subject.
Multi-Subject Models: Binary multi-subject classifiers performed best on average, effectively learning patterns across a set of users in a specific condition, or set of conditions. Multi-class classification also showed promise, surmounting random chance in all but one condition: Flappy Passive. Like the Flappy Passive condition, some multi-class models struggled to distinguish between adjacent classes of agent performance. Task-specific challenges and participant attention may have been an influence, as the Flappy Active multi-class model performed much better than its Passive counterpart. Although model performance was generally consistent across conditions, Lunar Lander and Flappy Bird conditions showed higher variance and lower performance compared to both Robot conditions, shown in Table 1.
Cross-Validation Studies: Cross-validation studies revealed little success across conditions and participants. Only the Robot Passive condition demonstrated any notable potential for zero-shot transferability when evaluated on users within the same condition ($F1=0.54$, $\sigma=0.01$). This finding aligns with prior work and suggests that developing generalizable BCI models, as explored in (huangfNIRS2MW2021), remains an open challenge. These findings motivate the application of domain adaptation and deep learning techniques to improve robustness across participants and tasks  (wang_taming_2021; eastmond_deep_2022).
@@
Regression Performance: Most regression models performed particularly well, highlighting the ability to extract deeper complexity from fNIRS data beyond scalar feedback. Figure 7 illustrates the performance of MLP, K-Nearest Neighbors and Random Forest regression models in the Robot Passive condition. Regressors trained on Active conditions performed marginally better on average ($R^{2}=0.81$) than Passive conditions ($R^{2}=0.77$). However, these models were unable to generalize to new participants. Finetuning with 30% of the target participant’s data increase these scores, but did not raise performance to a sufficient degree for application for any condition other than Lunar Active condition (Cross-Sub $R^{2}=0.01$, Fine-Tuned $R^{2}=0.23$). Regression models may need more target data to calibrate multi-subject models.
NASA-TLX Results: NASA-TLX allows participants to self-report mental workload through six different lenses: Mental, Temporal, Performance, Frustration, Physical and overall Effort Demands. Self-reported cognitive load during Active tasks was higher for nearly every category. Lunar Lander Active tasks were particularly high across all categories. Passive tasks were least demanding, and participants verbally described watching the autonomous Lunar Lander and Flappy Bird agent as ”boring”. These experiences may contribute to slightly lower Passive condition performance.
@@ Discussion
Our results indicate that both binary and multi-class classifiers are capable of identifying performance categories with statistically meaningful performance, and regression analysis further reveals that neural responses contain fine-grained evaluative signals beyond categorical labels. Our cross-participant evaluations highlight potential promise for generalization in the Robot Passive condition, suggesting that task structure and participant engagement may play critical roles in model transferability.
@@
A natural next step would be to integrate these models into real-time RLHF frameworks to enable adaptive agent behavior based on implicit neural feedback. Future work should also explore personalized calibration strategies, and taking advantage of multi-modal feedback (e.g., EEG, EMG, GSR) for greater insights into internal human assessments. Finally, we highlight the need to expand datasets to further support cross-subject generalization.
@@ Conclusion
This paper demonstrates a measurable relationship between fNIRS data and agent performance, showing that fNIRS signals can be decoded into meaningful evaluative feedback. We trained models to distinguish levels of agent performance based solely on passive neural responses, and repeated our procedures with active demonstration to serve as a benchmark against explicit feedback modalities. We show that passive fNIRS signals can be mapped to various levels of agent performance, laying the groundwork for enabling seamless integration of neural feedback into RLHF systems.

### v3

@@ Results
Single-Subject Models: Single-subject models successfully distinguished between binary ($F1=0.79$) and multi-class ($F1=0.75$) levels of agent performance. Multi-class classifiers successfully differentiated between optimal, suboptimal, and worst-case behavior with little variance. However, cross-validation studies revealed that no single-subject model was able to predict agent performance from the neural data of another subject.
Multi-Subject Models: Binary multi-subject classifiers performed best on average, effectively learning patterns across a set of users in a specific condition, or set of conditions. Multi-class classification also showed promise, exceeding random chance in all but one condition: Flappy Passive. Like the Flappy Passive condition, some multi-class models struggled to distinguish between adjacent classes of agent performance. Task-specific challenges and variations in participant attention may have influenced performance, as the Flappy Active multi-class model performed much better than its Passive counterpart. Although model performance was generally consistent across conditions, Lunar Lander and Flappy Bird conditions showed higher variance and lower performance compared to both Robot conditions, shown in Table 1.
Cross-Validation Studies: Cross-validation studies revealed little success across conditions and participants. Only the Robot Passive condition demonstrated any notable potential for zero-shot transferability when evaluated on users within the same condition ($F1=0.54$, $\sigma=0.01$). This finding aligns with prior work and suggests that developing generalizable BCI models, as explored in (Huang et al. 2021), remains an open challenge. These findings motivate the application of domain adaptation and deep learning techniques to improve robustness across participants and tasks (Wang et al. 2021; Eastmond et al. 2022).
@@
Regression Performance: Most regression models performed particularly well, highlighting the ability to extract deeper complexity from fNIRS data beyond scalar feedback. Figure 7 illustrates the performance of MLP, K-Nearest Neighbors, and Random Forest regression models in the Robot Passive condition. Regressors trained on Active conditions performed marginally better on average ($R^{2}=0.81$) than Passive conditions ($R^{2}=0.77$). However, these models were unable to generalize to new participants. Finetuning with 30% of the target participant’s data increased these scores, but did not raise performance to a sufficient degree for application in any condition other than Lunar Active (Cross-Sub $R^{2}=0.01$, Fine-tuned $R^{2}=0.23$). Regression models may need more target data to calibrate multi-subject models.
NASA-TLX Results: NASA-TLX allows participants to self-report mental workload through six different lenses: Mental, Temporal, Performance, Frustration, Physical, and overall Effort Demands. Self-reported cognitive load during Active tasks was higher for nearly every category (Figure 8). Lunar Lander Active tasks were particularly high across all categories. Passive tasks were least demanding; participants verbally described watching the autonomous Lunar Lander and Flappy Bird agent as “boring”. These experiences may contribute to slightly lower Passive condition performance.
@@ Discussion
Our results indicate that both binary and multi-class classifiers are capable of identifying performance categories with meaningful accuracy, and regression analysis further reveals that neural responses contain fine-grained evaluative signals beyond categorical labels. Our cross-subject evaluations highlight promising potential for generalization in the Robot Passive condition, suggesting that task structure and participant engagement may play critical roles in model transferability.
@@
A natural next step would be to integrate these models into real-time RLHF frameworks to enable adaptive agent behavior based on implicit neural feedback. Future work should also explore personalized calibration strategies, and take advantage of multi-modal feedback (e.g., EEG, EMG, GSR) for greater insights into internal human assessments. Finally, we highlight the need to expand datasets to further support cross-subject generalization.
@@ Conclusion
This paper demonstrates a measurable relationship between fNIRS data and agent performance, showing that fNIRS signals can be decoded into meaningful evaluative feedback. We trained models to distinguish levels of agent performance based solely on passive neural responses, and repeated our procedures using active demonstration to serve as a benchmark against explicit feedback modalities. We show that passive fNIRS signals can be mapped to various levels of agent performance, laying the groundwork for enabling seamless integration of neural feedback into RLHF systems.